OpenStudy (anonymous):

simplify and state the excluded values for each variable #1 3x^2-10x+3/3x^2-7x+2 #2 z^2+z-12/z^2-z-12

OpenStudy (anonymous):

\[\frac{3x^2-10x+3}{3x^2-7x+2}\]?

OpenStudy (anonymous):

wait up the first one is the exact same thing i just wrote isnt ot?

OpenStudy (anonymous):

qstion is fine

OpenStudy (anonymous):

simplify it to get (3x-1)(x-3)/(3x-1)(x-2) = x-3/x-2

OpenStudy (anonymous):

excluded value x=2

OpenStudy (anonymous):

\[\frac{(3x-1)(x-3)}{(3x-1)(x-2)}=\frac{x-3}{x-2}\]

OpenStudy (anonymous):

x=2 you need number 2 ?

OpenStudy (anonymous):

the excluded value for x is 2 coz 2-2 =0 which makes the fraction infinite

OpenStudy (anonymous):

next one answer is (z+4)(z-3) / (z-4)(z+3)

OpenStudy (anonymous):

in this excluded values for z are -3, 4

OpenStudy (anonymous):

Yeah nancy I need help with #2

OpenStudy (anonymous):

roopadr give u answer